Draw the shear and moment diagrams for the beam

250 lb/ft B 150 lb-ft 150 lb-ft 20 ft V, 1b
Answer #1

Jio Page No.: Х son Isolatt 150-euft 150 lbst А he 13 X 12 oft RB Х Moment about A = 150 + ( (-250 x20x10 -150 + 20 R = 0 .:Jio Page No.: shear force diagram 2500.16 A B - 2500 lb Ben ending Moment diagram 12350 lb-ft 1 - 70 = 20 X 10 - 150 lb-ft d